# and extensions of the log-Sobolev and Talagrand inequalities

### 04 Mar 2024 v1

@@ B.1 Proof of Proposition 6
$\mathinner{\!\left\lVert\theta_{t}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert X_{t}\right\rVert}^{2}\right]\leq 2\left(\mathinner{\!\left\lVert\theta_{\dagger}\right\rVert}^{2}+\mathinner{\!\left\lVert x_{\dagger}\right\rVert}^{2}+\mathinner{\!\left\lVert\theta_{0}-\theta_{\dagger}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert X_{0}-x_{\dagger}\right\rVert}^{2}\right]+\frac{d_{x}}{\lambda}\right)\quad\forall t\geq 0.$
@@
applying Itô’s formula [30, Theorem 4.2.1] and noting that $X_{t}$ has law $q_{t}$, we find
$\operatorname{d\!}\xi_{t}=\left[2\left\langle{(\theta_{t}-\theta_{\dagger},X_{t}-x_{\dagger})},{(\mathbb{E}\left[\nabla_{\theta}\ell(\theta_{t},X_{t})\right],\nabla_{x}\ell(\theta_{t},X_{t}))}\right\rangle+d_{x}\right]\operatorname{d\!}t+\sqrt{2}\left\langle{X_{t}-x_{\dagger}},{\operatorname{d\!}W_{t}}\right\rangle.$ (43)
@@
In turn, taking expectations of (43) yields $\operatorname{d\!}\mathbb{E}\left[\xi_{t}\right]\leq[-\lambda\mathbb{E}\left[\xi_{t}\right]+d_{x}]\operatorname{d\!}t\enskip\forall t\geq 0$ and
$\mathbb{E}\left[\xi_{t}\right]\leq e^{-\lambda t}\left(\mathbb{E}\left[\xi_{0}\right]-\frac{d_{x}}{\lambda}\right)+\frac{d_{x}}{\lambda}\leq\max\bigg{(}\mathbb{E}\left[\xi_{0}\right],\frac{d_{x}}{\lambda}\bigg{)}\leq\mathbb{E}\left[\xi_{0}\right]+\frac{d_{x}}{\lambda}\quad\forall t\geq 0,$
@@
Proof of Proposition 6.
This is an adaptation of the arguments in [7, Theorem 1.7], [8, Proposition 1], and [24, Proposition 3.1]. Fix a $T>0$ and define the map
$\Psi\mathrel{\mathop{:}}\mathcal{C}([0,T],\mathbb{R}^{d_{\theta}})\times\mathcal{P}_{2}(\mathcal{C}([0,T],\mathbb{R}^{d_{x}}))\mapsto\mathcal{C}([0,T],\mathbb{R}^{d_{\theta}})\times\mathcal{P}_{2}(\mathcal{C}([0,T],\mathbb{R}^{d_{x}})),$
$(\theta,\nu)\mapsto\Psi(\theta,\nu)$, associating to each pair $(\theta,\nu)$ the law on path space of the solution $(\theta_{t},X_{t})_{t\leq T}$ of the SDE (6), which is convenient to rewrite as
$\operatorname{d\!}(\theta_{t},X_{t})=b(\theta_{t},X_{t},\nu_{t})\operatorname{d\!}t+\sigma\operatorname{d\!}W_{t}$
where $b(\theta_{t},X_{t},\nu_{t})\mathrel{\mathop{:}}=[\int\nabla_{\theta}\ell(\theta_{t},x)\nu_{t}(\operatorname{d\!}x),\nabla_{x}\ell(\theta_{t},X_{t})]^{\intercal}$, $\sigma\mathrel{\mathop{:}}=\sqrt{2}[0_{d_{\theta}},1_{d_{x}}]^{\intercal}$ and $\nu_{t}$ is the time marginal of $\nu$ at $t$. Notice that $\Psi$ is well defined by Proposition 19. We aim to show that $\Psi$ has a unique fixed point, which is both necessary and sufficient for $(\theta_{t},X_{t})_{t\leq T}$ to solve the SDE above in the strong sense. To do that, we first define the straightforward extension of the metric $\mathsf{d}$ to the path space $\mathcal{C}([0,T],\mathbb{R}^{d_{\theta}})\times\mathcal{P}_{2}(\mathcal{C}([0,T],\mathbb{R}^{d_{x}}))$:
$\mathsf{d}_{T}((\theta_{[0,T]},\nu_{[0,T]}),(\theta_{[0,T]}^{\prime},\nu_{[0,T]}^{\prime}))\mathrel{\mathop{:}}=\sup_{t\leq T}\sqrt{\mathinner{\!\left\lVert\theta_{t}-\theta_{t}^{\prime}\right\rVert}^{2}+W_{2}(\nu_{t},\nu_{t}^{\prime})}.$
Let $(\theta_{t},X_{t})_{t\leq T},(\theta_{t}^{\prime},X_{t}^{\prime})_{t\leq T}$ be two solutions of (6) satisfying $\theta_{0}=\theta_{0}^{\prime}$ and $X_{0}=X_{0}^{\prime}$. For some rolling constant $c$ we compute by Jensen’s inequality and then Lipschitz continuity and the Kantorovich–Rubinstein duality formula,
$\mathbb{E}\left[\mathinner{\!\left\lVert\theta_{t}-\theta_{t}^{\prime}\right\rVert}^{2}+\mathinner{\!\left\lVert X_{t}-X_{t}^{\prime}\right\rVert}^{2}\right]$ $\leq\mathbb{E}\left[\mathinner{\!\left\lVert\int_{0}^{t}(b(\theta_{s},X_{s},\nu_{s})-b(\theta_{s}^{\prime},X_{s}^{\prime},\nu_{s}^{\prime}))\operatorname{d\!}s\right\rVert}^{2}\right]$
$\leq T\int_{0}^{t}\mathbb{E}\left[\mathinner{\!\left\lVert b(\theta_{s},X_{s},\nu_{s})-b(\theta_{s}^{\prime},X_{s}^{\prime},\nu_{s}^{\prime})\right\rVert}^{2}\operatorname{d\!}s\right]$
$\leq cT\int_{0}^{t}\mathbb{E}\left[\mathinner{\!\left\lVert\theta_{s}-\theta_{s}^{\prime}\right\rVert}^{2}+\mathinner{\!\left\lVert X_{s}-X_{s}^{\prime}\right\rVert}^{2}+W_{1}(\nu_{s},\nu_{s}^{\prime})^{2}\right]\operatorname{d\!}s$
Next, using the fact $W_{1}\leq W_{2}$,
$\int_{0}^{t}W_{1}(\nu_{r},\nu_{r}^{\prime})^{2}\operatorname{d\!}s\leq\int_{0}^{t}\sup_{r\leq s}W_{1}(\nu_{r},\nu_{r}^{\prime})^{2}\operatorname{d\!}s\leq\int_{0}^{T}\mathsf{d}_{s}((\theta_{[0,s]},\nu_{[0,s]}),(\theta_{[0,s]}^{\prime},\nu_{[0,s]}^{\prime}))^{2}\operatorname{d\!}s.$
Combining these two estimates, Grönwall’s inequality then shows that for all $t\leq T$,
$\mathbb{E}\left[\mathinner{\!\left\lVert\theta_{t}-\theta_{t}^{\prime}\right\rVert}^{2}+\mathinner{\!\left\lVert X_{t}-X_{t}^{\prime}\right\rVert}^{2}\right]\leq c_{T}\int_{0}^{T}\mathsf{d}_{s}((\theta_{[0,s]},\nu_{[0,s]}),(\theta_{[0,s]}^{\prime},\nu_{[0,s]}^{\prime}))^{2}\operatorname{d\!}s.$
with $c_{T}\mathrel{\mathop{:}}=cTe^{T}$. Using the definition of $\Psi$ and $\mathsf{d}_{T}$, taking the supremum across $t\leq T$,
$\mathsf{d}_{T}(\Psi(\theta_{[0,T]},\nu_{[0,T]}),\Psi(\theta_{[0,T]}^{\prime},\nu_{[0,T]}^{\prime}))^{2}\leq c_{T}\int_{0}^{T}\mathsf{d}_{s}((\theta_{[0,s]},\nu_{[0,s]}),(\theta_{[0,s]}^{\prime},\nu_{[0,s]}^{\prime}))^{2}\operatorname{d\!}s.$
Let $\Psi^{k}$ denote the $k$-fold composition of $\Psi$. Iterating the inequality above $k$-times yields
$\mathsf{d}_{T}(\Psi^{k}(\theta_{[0,T]},\nu_{[0,T]}),\Psi^{k}(\theta_{[0,T]}^{\prime},\nu_{[0,T]}^{\prime}))^{2}$ $\leq(c_{T})^{k}\int_{0}^{T}\frac{s^{k}}{(k-1)!}\mathsf{d}_{s}((\theta_{[0,s]},\nu_{[0,s]}),(\theta_{[0,s]}^{\prime},\nu_{[0,s]}^{\prime}))^{2}\operatorname{d\!}s$
$\leq\frac{(c_{T})^{k}}{k!}\mathsf{d}_{T}((\theta_{[0,T]},\nu_{[0,T]}),(\theta_{[0,T]}^{\prime},\nu_{[0,T]}^{\prime}))^{2}.$
Letting $k\rightarrow\infty$ shows that $\Psi^{k}$ is a contraction for some $k\in\mathbb{N}$ and the first claim follows by the fixed point theorem. Now we want to show that the law of the solution to the SDE (6) solves the gradient flow (5) weakly, in the sense of Assumption 2. We adapt an argument in [30, Section 8.3]. Let $\phi$ in $\mathcal{C}_{c}^{2}(\mathbb{R}^{d_{x}})$, where $\mathcal{C}_{c}^{2}(\mathbb{R}^{d_{x}})$ is the space of twice continuously differentiable compactly supported functions on $\mathbb{R}^{d_{x}}$. Formula (7.3.1) in [30] shows
Now consider the system
$\dot{\theta}_{t}=\int\nabla_{\theta}\ell(\theta_{t},x)\nu_{t}(\operatorname{d\!}x),\quad M_{t}\mathrel{\mathop{:}}=\phi(X_{t})-\int_{0}^{t}\bigg{(}\left\langle{\nabla_{x}\ell(\theta_{s},X_{s})},{\nabla_{x}\phi(X_{s})}\right\rangle+\Delta_{x}\phi(X_{s})\bigg{)}\operatorname{d\!}s$
with $\nu_{t}=\text{Law}(X_{t})$. Comparing with the expression for $\phi(X_{t})$ above we notice that $M_{t}$ is a martingale with respect to the natural filtration generated by $(\theta_{t},X_{t})_{t\geq 0}$, as it corresponds to the Itô integral of an adapted, square integrable process against Brownian motion. Taking expectations and the time derivative in $M_{t}$ above and then integrating by parts shows that
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\phi(x)\nu_{t}(\operatorname{d\!}x)$ $=-\int\left\langle{\nabla_{x}\ell(\theta_{t},x)},{\nabla_{x}\phi(x)}\right\rangle\nu_{t}(\operatorname{d\!}x)+\int\Delta_{x}\phi(x)\nu_{t}(\operatorname{d\!}x)\quad\forall\phi\in\mathcal{C}_{c}^{2}(\mathbb{R}^{d_{x}}).$
B.2 Proof of Lemma 10
$=\mathinner{\!\left\lVert\theta_{t}-\theta_{*}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert X_{t}^{1}-X_{*}^{1}\right\rVert}^{2}\right]=\mathsf{d}((\theta_{t},q_{t})),(\theta_{*},\pi_{\theta_{*}}))^{2}$
The result follows via Theorem 5 and Proposition 6. ∎
B.3 Proof of Lemma 11
@@
$\left|\mathbb{E}\left[G_{t}^{N}\right]\right|\leq L\sqrt{\frac{2\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]}{N}\left(B_{0}+\frac{d_{x}}{\lambda}\right)}.$ (47)
Because $(\operatorname{d\!}/\operatorname{d\!}t)\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}=2^{-1}\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{-1/2}(\operatorname{d\!}/\operatorname{d\!}t)\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]$, (46,47) imply that
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}\leq-\lambda\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}+L\sqrt{\frac{2}{N}\left(B_{0}+\frac{d_{x}}{\lambda}\right)}.$
@@
$\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}$ $\leq e^{-\lambda t}\sqrt{\bar{\xi}_{0}^{N}}+\frac{(1-e^{-\lambda t})L}{\lambda}\sqrt{\frac{2}{N}\left(\mu_{0}+\frac{d_{x}}{\lambda}\right)},$
@@ B.4 Proof of Lemma 12
The following proof extends the arguments in [16, Lemma S2] or [11, Chapter 4.1]. We introduce the linear interpolation of the Euler-Maruyama discretization of (32):
@@
Furthermore, under Assumptions 4 and 5 we have, by [29, Theorem 2.1.12], there holds the co-coercivity property
@@
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\mathinner{\!\left\lVert\nabla\ell(\bar{Y}_{s}^{n})-\nabla\ell(\bar{Y}_{kh}^{n})\right\rVert}^{2}\right]\leq 220hL^{2}(L^{2}h(B_{0}+d_{x}/\lambda+d_{x}).$ (53)
@@
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\mathinner{\!\left\lVert X_{t}^{n}\right\rVert}^{2}+\mathinner{\!\left\lVert\Theta_{t}^{N}\right\rVert}^{2}\right]\leq 2\left(\mathinner{\!\left\lVert\theta_{0}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert\bar{X}_{0}\right\rVert}^{2}\right]+\frac{d_{x}}{\lambda}\right)$
@@
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathbb{E}\left[\xi_{t}^{N}\right]=-2\mathbb{E}\left[\frac{1}{N}\sum_{n=1}^{N}\left\langle{(X^{n}_{t},\Theta_{t}^{N})},{\nabla\ell(\Theta_{t}^{N},X^{n}_{t})}\right\rangle\right]+2d_{x}\leq-2\lambda\mathbb{E}\left[\xi_{t}^{N}\right]+2d_{x},$

### 11 Apr 2024 v2

@@ B.3 Proof of Lemma 11
$\mathinner{\!\left\lVert\theta_{t}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert X_{t}\right\rVert}^{2}\right]\leq 2\left(\mathinner{\!\left\lVert\theta_{\dagger}\right\rVert}^{2}+\mathinner{\!\left\lVert x_{\dagger}\right\rVert}^{2}+\mathinner{\!\left\lVert\theta_{0}-\theta_{\dagger}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert X_{0}-x_{\dagger}\right\rVert}^{2}\right]+\frac{2d_{x}}{\lambda}\right)\quad\forall t\geq 0.$
If $(\theta_{\dagger},x_{\dagger})=(0,0)$, the above also holds without the factor of 2 in the right hand side.
@@
applying Itô’s formula [35, Theorem 4.2.1] and noting that $X_{t}$ has law $q_{t}$, we find
$\operatorname{d\!}\xi_{t}=2\left[\left\langle{(\theta_{t}-\theta_{\dagger},X_{t}-x_{\dagger})},{(\mathbb{E}\left[\nabla_{\theta}\ell(\theta_{t},X_{t})\right],\nabla_{x}\ell(\theta_{t},X_{t}))}\right\rangle+d_{x}\right]\operatorname{d\!}t+2^{3/2}\left\langle{X_{t}-x_{\dagger}},{\operatorname{d\!}W_{t}}\right\rangle.$ (47)
@@
In turn, taking expectations of (47) yields $\operatorname{d\!}\mathbb{E}\left[\xi_{t}\right]\leq[-\lambda\mathbb{E}\left[\xi_{t}\right]+2d_{x}]\operatorname{d\!}t\enskip\forall t\geq 0$ and
$\mathbb{E}\left[\xi_{t}\right]\leq e^{-\lambda t}\left(\mathbb{E}\left[\xi_{0}\right]-\frac{2d_{x}}{\lambda}\right)+\frac{2d_{x}}{\lambda}\leq\max\bigg{(}\mathbb{E}\left[\xi_{0}\right],\frac{2d_{x}}{\lambda}\bigg{)}\leq\mathbb{E}\left[\xi_{0}\right]+\frac{2d_{x}}{\lambda}\quad\forall t\geq 0,$
@@
Proof of Lemma 11.
@@
$\left|\mathbb{E}\left[G_{t}^{N}\right]\right|\leq L\sqrt{\frac{2\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]}{N}\left(B_{0}+\frac{2d_{x}}{\lambda}\right)}.$ (51)
Because $(\operatorname{d\!}/\operatorname{d\!}t)\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}=2^{-1}\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{-1/2}(\operatorname{d\!}/\operatorname{d\!}t)\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]$, (50,51) imply that
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}\leq-\lambda\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}+L\sqrt{\frac{2}{N}\left(B_{0}+\frac{2d_{x}}{\lambda}\right)}.$
@@
$\mathbb{E}\left[\bar{\xi}_{t}^{N}\right]^{1/2}$ $\leq e^{-\lambda t}\sqrt{\bar{\xi}_{0}^{N}}+\frac{(1-e^{-\lambda t})L}{\lambda}\sqrt{\frac{2}{N}\left(B_{0}+\frac{2d_{x}}{\lambda}\right)},$
@@ B.4 Proof of Lemma 12
The following proof extends the arguments in [18, Lemma S2] or [13, Chapter 4.1]. We introduce the linear interpolation of the Euler-Maruyama discretization of (32):
@@
Furthermore, under Assumptions 3 and 4 we have, by [34, Theorem 2.1.12], there holds the co-coercivity property
@@
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\mathinner{\!\left\lVert\nabla\ell(\bar{Y}_{s}^{n})-\nabla\ell(\bar{Y}_{kh}^{n})\right\rVert}^{2}\right]\leq 220hL^{2}(L^{2}h(B_{0}+2d_{x}/\lambda+d_{x}).$ (57)
@@
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\mathinner{\!\left\lVert X_{t}^{n}\right\rVert}^{2}+\mathinner{\!\left\lVert\Theta_{t}^{N}\right\rVert}^{2}\right]\leq\left(\mathinner{\!\left\lVert\theta_{0}\right\rVert}^{2}+\mathbb{E}\left[\mathinner{\!\left\lVert\bar{X}_{0}\right\rVert}^{2}\right]+\frac{2d_{x}}{\lambda}\right)$
@@
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathbb{E}\left[\xi_{t}^{N}\right]=-2\mathbb{E}\left[\frac{1}{N}\sum_{n=1}^{N}\left\langle{(X^{n}_{t},\Theta_{t}^{N})},{\nabla\ell(\Theta_{t}^{N},X^{n}_{t})}\right\rangle\right]+2d_{x}\leq-\lambda\mathbb{E}\left[\xi_{t}^{N}\right]+2d_{x},$